A person travels by car from one city to another. She drives for 28.4 min @ 67.2 km/h, 12.3 min. @ 117km/h, 38.7 min @ 36.2km/h and spends 15.1 min along the way eating lunch and buying gas. Determine the distance between the cities along the route. Ans in km
1
Expert's answer
2011-02-01T05:46:37-0500
S = V1T1 + V2T2 + V3T3 +V4T4 =
=28.4/60 [h] * 67.2 [km/h] + 12.3/60[h]* 117[km/h] + 38.7/60[h] * 36.2[km/h] 15.1/60[h] *0[km/h]= 79.142 [km]


Answer: 79.142 km.